Equivalent Units of Production and Related Costs The charges to Work in Process—Baking Department for a...

  1. Equivalent Units of Production and Related Costs

    The charges to Work in Process—Baking Department for a period as well as information concerning production are as follows. The Baking Department uses the weighted average method, and all direct materials are placed in process during production.

    Work in Process—Baking Department
    Bal., 900 units, 40% completed 2,466 To Finished Goods, 8,100 units ?
    Direct materials, 8,400 units 34,500
    Direct labor 16,200
    Factory overhead 8,574
    Bal., 1,200 units, 60% completed ?

    a. Determine the number of whole units to be accounted for and to be assigned costs.
    units

    b. Determine the number of equivalent units of production.
    units

    c. Determine the cost per equivalent unit.
    $ per equivalent unit

    d. Determine the cost of units transferred to Finished Goods.
    $

    e. Determine the cost of units in ending Work in Process.
    $
Answer #1

a) Total Units accounted for = 900+8400 = 9300

b) Equivalent unit = 8100+(1200*60%) = 8820

c) Cost per equivalent unit = (2466+34500+16200+8574)/8820 = 7

d) Cost of unit transferred = 7*8100 = 56700

e) Cost of ending WIP = 720*7 = 5040